OpenStudy (anonymous):

Help Please A right triangle has base (x - 6) units and height (6x - 12) units. Part A: What is the square of the length of the hypotenuse of the triangle? Show your work. Part B: What is the area of the triangle? Part C: Using the solution obtained in Part B, explain the closure property of multiplication of polynomials.

OpenStudy (anonymous):

im guessing you have to square (x-6)^2 and (6x-12)^2

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

and then for part b i solve it right

OpenStudy (anonymous):

tell me if im wrong please

OpenStudy (anonymous):

anyone

OpenStudy (anonymous):

know

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Part A wants the length of the hypotenuse in squared form. Do you know how to find the length of the hypotenuse when you know the length of the other two sides?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry

OpenStudy (imstuck):

Yes, you're right. For part B, the answer I get it 6 *sqrt (x-2)^2 + (x-6)^2 = c

OpenStudy (anonymous):

 \[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

We need c^2

OpenStudy (anonymous):

ok how do i do that

OpenStudy (anonymous):

Ok, well in this case we know the value of the base and the height, so it would look like this:

OpenStudy (anonymous):



OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

i dont see anything

OpenStudy (anonymous):

Ok, do you know what a hypotenuse of a triangle is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

It's the longest side of a right triangle. It's also a right triangle because it has a 90 degree angle.

OpenStudy (anonymous):

ok so do i have to take the height and square it

OpenStudy (anonymous):

so (6x-12)^2

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] We have to square the height, and add it to the square of the base.

OpenStudy (anonymous):

so (x-6)^2+(6x-12)^2

OpenStudy (anonymous):

Yep, now when you simplify that don't remove the square.

OpenStudy (anonymous):

ok so how do i simplify

OpenStudy (anonymous):

sorry ive been testing for 5 hours my brains kind of fried

OpenStudy (anonymous):

so do i do the foil method

OpenStudy (anonymous):

You would use the foil method if this was a quadratic we were going to simplify, but that's actually completely unrelated

OpenStudy (anonymous):

\[(x - 6)^2 + (6x - 12)^2 = c^2\] First we should simplify the x values

OpenStudy (anonymous):

What is x + 6x?

OpenStudy (anonymous):

6x^2

OpenStudy (anonymous):

That would be 6x * x

OpenStudy (anonymous):

\[x + x = 2x\] \[x * x = x^2\]

OpenStudy (anonymous):

oh then im not sure sorry

OpenStudy (anonymous):

is it 7x or 8x

OpenStudy (anonymous):

for x + 6x, remember that there's actually a hidden 1 beside the x. 1x + 6x = 7x

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

Now -6 - 12 simplified is

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

so for part A: \[c = (7x - 18)^2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so what do i do for b

OpenStudy (anonymous):

For part B we need the area of the triangle The formula for the area of a right triangle is: \[\frac{ ab }{ 2 }\]

OpenStudy (anonymous):

And again: A = x - 6 B = 6x - 12

OpenStudy (anonymous):

ok so i have to solve for x

OpenStudy (anonymous):

Yep, I think I can see some foiling up ahead

OpenStudy (anonymous):

ok so what do i have to foil

OpenStudy (anonymous):

oh wait never mind its A = x - 6 B = 6x - 12 that i have to foil right

OpenStudy (anonymous):

\[Area = \frac{ ab }{ 2 }\] \[Area =\frac{ (x - 6)(6x - 12) }{ 2 }\]

OpenStudy (anonymous):

The binomial on top should be turned into a quadratic I believe

OpenStudy (anonymous):

ok so would be

OpenStudy (anonymous):

im trying to draw it

OpenStudy (anonymous):

so the second part would be (3x-6)

OpenStudy (anonymous):

but for the first it would be (?-3) im not sure what x/2 is

OpenStudy (anonymous):

Firs just foil (x−6)(6x−12)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so the same as before

OpenStudy (anonymous):

(7x-18)just not squared this time

OpenStudy (anonymous):

The other time we added them together, this time we're multiplying. Do you know the steps to foiling?

OpenStudy (anonymous):

oh yeah i understand

OpenStudy (anonymous):

so

OpenStudy (anonymous):

(6x^2 am i starting out right

OpenStudy (anonymous):

That was also my first term

OpenStudy (anonymous):

ok so for the rest is it x times 12 or -12

OpenStudy (anonymous):

For the outers it's x * (-12)

OpenStudy (anonymous):

ok so what would that be im always confused when it comes to multiplying x

OpenStudy (anonymous):

what it be -12x

OpenStudy (anonymous):

-12x would be correct

OpenStudy (anonymous):

So now should be the inners

OpenStudy (anonymous):

ok so (6x^2-(-12)9-36-72)

OpenStudy (anonymous):

oops i made an error

OpenStudy (anonymous):

(6x^2-(-12)(-36-72)

OpenStudy (anonymous):

When simplified I got:\[6x^2 - 48x + 72\]

OpenStudy (anonymous):

i so i did it wrong where did i mess up

OpenStudy (anonymous):

I think there was a problem with the positives and negatives of 12 and 72

OpenStudy (anonymous):

oh ok so the final answer is 3x^2-42x+36

OpenStudy (anonymous):

i mean 24

OpenStudy (anonymous):

3x^2-24x+36

OpenStudy (anonymous):

Yep, so I'll just re-write it:

OpenStudy (anonymous):

\[3x^2 - 24x + 36\]

OpenStudy (anonymous):

First of all do you recognize what type of expression this is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

This is a quadratic because the highest power is 2